# Expertized Caption Auto-Enhancement for Video-Text Retrieval

### v1

@@ 4.1 Datasets
Table 2: Comparison with SOTA methods on MSVD and DiDeMo datasets (underline marks the best ViT-based results)
4.2 Implementation Details
@@
Following the standard text preprocessing [20], each sentence in original texts and video-derived captions is split into word tokens with a max length of 70 using a CLIP tokenizer. For videos, we sample 8 frames from each item as visual data.
@@
This paper reports the results of our method on multiple backbones, i.e., ViT-B/16, ViT-L/14, ViT-H/14, ViT-G/14, compared to SOTA video-text retrieval methods, showing the effect of model scale on performance as well. Ten captions are derived for each video. Sixteen experts are learned in the ECS module, two of which are activated to select appropriate representations in both training and inference processes. In the training phase, the CSI module is pre-learned to acquire the optimal prompt for caption generation, while the ECS module is co-trained alongside the backbone parameters. In the evaluation phase, multiple captions are derived from each video using the optimal prompt, together with the video itself as video-side expressions. (See Supplemental Materials for the learned best prompt and examples of video-derived captions.) The Captioner and Prompt engineer are implemented by GPT-4o, a large cross-modal inference model recently released by OpenAI in 2024. Besides the comparison experiment, we encapsulate ExCae as a plug-in unit and test its effect when introduced into existing video-text retrieval methods; specifically, we enhance video expression via CSI and professionalize the video-text matching process by ECS. Ablation studies and analytical experiments are conducted on MSR-VTT using ViT-G/14 as the base model.
@@
Our model is implemented in PyTorch and trained on 8 NVIDIA A800 GPUs in a batch size of 16. We initialize the text and video encoder with pretrained weights from CLIP [20] and use the Adamax optimizer with a learning rate of 4e-6. Experimental results are evaluated by standard video-text retrieval metrics: Recall at Rank K (R@K, higher is better) and Mean Rank (MeanR, lower is better). R@K denotes the percentage of correct samples in the top K retrieved samples, where K is set to 1, 5 and 10, the same as in the previous works [31, 9]. MeanR records the average rank of correct items in the retrieved list.
4.3 Comparision to State-of-the-arts
@@
Results of our method are compared to those of the SOTAs, including pre-trained foundation models like InternVideo [24], UMT-L [11], mPLUG-2 [28], VAST [6], and ViT-based methods that adapt the image-text models to the video-text task. Among the ViT-based methods, CLIP2TV [8], DRL [23], Clip4Clip [15] and CLIP-VIP [31] require the acquisition of auxiliary data for post-pretraining and/or adaptation. In contrast, TS-Net [14], X-CLIP [16], DMAE [9] and Cap4Video [26] are learned without extra data.
@@
Results in Table 1 show that, on the MSR-VTT dataset, our method is superior not only to the ViT-based methods but also to the pre-training models relying on large-scale pre-training data. The results of existing methods show that the Top-1 recall accuracy could hardly transcend 50% without assistance from extra data. Instead, great improvement (up to 10%) is obtained when additional data is involved, either they are used for pre-training or fine-tuning. Despite not using any extra data, the Top-1 recall accuracy of our method (ExCae) boosts to 55.0% and 56.1% in text-to-video and video-to-text retrieval, respectively, when using ViT-B/16 as the base model. We are also encouraged by the observation that ExCae’s performance demonstrates an increase in correlation with the enlarged scale of the backbone model, resulting in an around 7% rise of R@1 from ViT-L/14 to Vit-G/14. Our results have comprehensively surpassed the current SOTA pre-training models, such as mPLUG-2 and VAST, when using the Vit-G/14 as the backbone. The rising trend of performance indicates the potential for further improvement of our method through scaling up the base model. Similar results are obtained in the experiments on MSVD and DiDeMo datasets, as recorded in Table 2. Our results are comparable to or even superior to those of pre-training models. Compared to existing ViT-based methods, ExCae shows a resounding victory, with Top-1 recall accuracy suppressing the best one by 7% on MSVD and 13% on DiDeMo. These results validate that our method is robust to cross-modal retrieval tasks for data collected from different fields.
@@
We also demonstrate some video-text retrieval cases to present the performance of our method intuitively. The top three retrieved videos, when given a caption, are presented in Figure 4. In these cases, our method successfully picks the corresponding ground truth videos in the first place. The second and third hit videos are also highly related to the sentences in terms of content. Besides, Figure 5 compares our video-to-text retrieval results to the “Base” obtained without any caption enhancement. It is shown that the ground truth texts normally take the first rank when using our method. Compared to the base results, our method preferentially selects samples with semantics and descriptive perspectives similar to the ground truths. For example, our method successfully recognizes keywords of “voices”, “critic”, and “guitar” that are missed by the Base model. This is owing to the ECS module that adaptively learns the cross-modal matching for specific samples, benefiting the recognition of key semantics during video-text retrieval.
@@
4.4 Plug-in Experiment
ExCae is encapsulated and supplemented with existing video-text retrieval methods, whose results are compared to those of the original. As shown in Table 3, our method successfully increases retrieval results of different video-text retrieval methods. The Top-1 recall accuracy is boosted by 4.8%, 9.4% and 7.7% for DRL, TSR-Net and Clip4Clip, respectively. It is observed that the results of plug-in experiments are similar to those obtained by our method with the same backbone (refer to Table 1). This indicates that ExCae has strong adaptability and expandability, which can flexibly combine with different models and constantly upgrade its performance as the technology evolves.
@@ 4.4 Plug-in Experiment
Table 3: Results of plug-in experiments on MSR-VTT dataset.
@@
Table 4: Ablation studies of critical components
4.5 Ablation Studies
Our method expands the expression on the video side; thus, this section respectively evaluates the key components, CSI and ECS modules, under different input data cases. To test the effectiveness of the CSI module, we replace the optimal prompt with the initial one to simulate the ablation of the CSI module. The removal of the ECS module is realized by directly encoding all the video-side data as visual embeddings. As shown in Table 4, the R@1 is around 50% when only using the original video as video-side data without adding any components; this result is close to the existing ViT-based methods. The ECS module shows only slight positive effects on the original video. Similar results can be found if we merely change the video-side data from the original videos to video-derived captions. An obvious performance boost occurs when the original videos and video-derived captions are input in combination, validating the indispensability of these two data. In addition, we find that both CSI and ECS modules facilitate cross-modal retrieval. Despite in the absence of the other one, our method achieves over 6% gain in Top-1 recall accuracy. The effectiveness of the CSI and ECS modules proves the superiority of captions’ auto-enhancement and multi-perspective cross-modal matching. The best results are obtained by our full method, where both the CSI and ECS modules play mutually reinforcing roles in promoting video-text understanding.
4.6 Analytical Experiments
a) Effect of video-derived caption number ($N_{c}$)
Figure 6 plots the video-text retrieval results when raising $N_{c}$ from 1 to 10. The performance is really poor if only one caption is derived, with a Top-1 recall accuracy of lower than 30% on text-to-video retrieval. This result is inferior to any existing methods, reflecting that insufficient caption enhancement would even damage cross-modal learning. The reason is that all generated captions are forcibly activated to match the text if too few caption samples are input to the ECS module. Consequently, the original text may align with a description from a completely different perspective. Results are much better when increasing the number of generated captions to 3, achieving good performance close to the existing pre-training models. The performance further improves with the increase of the caption number, and the upward trend becomes stable when $N_{c}$ is greater than 7. This experiment suggests an approximate range of $[7,10]$ for the selection of video-derived caption numbers.
b) Effect of activated expert number ($N_{e}$)
Video-text retrieval results with different $N_{e}$ values are displayed in Figure 7. We also show the results obtained without the ECS module, labeled as 0 activated expert, as a reference. The retrieval results gradually improve as $N_{e}$ increases from 0 to 2. However, no further improvement in performance is achieved with the subsequent increase of activated experts. This may be because ECS would automatically favor partial experts and suppress the attention to less relevant expressions, although the majority of experts are activated.
@@ 4.6 Analytical Experiments
5 Discussion
5.1 Superiority
ExCae attains remarkable video-text retrieval performance primarily due to its efficient narrowing of the video-text modality gap. As shown in Figure 8 (a), the representations of video-text pairs are drawn closer to each other through self-learning diffusional data augmentation (CSI) and self-adaptive cross-modal matching (ECS). The modality gap $||\Delta||$ [12] is reduced from 0.82 to 0.75, from 0.79 to 0.67, and from 0.84 to 0.74 on MSR-VTT, MSVD and DiDeMo, respectively. Moreover, ExCae has advantages in objectivity, generalization and self-adaptation. 1) ExCae is a data-driven method, in which the query prompts and caption selection are learned according to the training data. No additional human work (despite initialization), such as prompt design and lexicon establishment, is involved in the whole learning process. Thus, ExCae effectively avoids the adverse effects of empiricism. 2) As the Captioner is asked to generate multi-angle captions using a diversity constraint, data with more expression perspectives are provided for model learning. Therefore, ExCae inherits the strength of generalization improvement from data augmentation. 3) ExCae also has personalized ability owing to the parametric learning of the ECS module. The router in the ECS module would adaptively select the best expressions for cross-modal retrieval, thus maintaining the robustness of ExCae in different applications.
@@ 5.1 Superiority
5.2 Convergence
Since the generated captions are changeless once the optimal prompt is determined, the cross-modal post-training of our method owns the same convergence as existing ViT-based methods. Regarding learning the CSI module, the LLM would optimize query prompts in the given direction. With the increase of iteration, the semantic changes in query prompts become smaller, and thus the caption evaluation scores tend to be stable (See Figure 8 (b)). According to expertise from experiments, the CSI module tends to converge in 400 iterations.
5.3 Storage and Efficiency
Although a prompt optimization process is additionally involved, we believe that the storage and computation cost of ExCae is appreciable, especially compared to the pre-training methods. With regard to storage, no extra data or related lexicon needs to be stored. We only need to occupy a very small amount of storage space to record several generated captions in each epoch, which greatly reduces the storage cost. For the computation cost, our method is more efficient than the multi-modal foundation models without a pre-training process. In comparison to the ViT-based methods, the only additional computation is from the training of the CSI module. However, prompt optimization is a one-off training process, and we have also found that the learned prompt has generalization ability across different data sets. Therefore, the efficiency of our approach is comparable to existing video-text retrieval methods.
6 Conclusion
This paper proposes strengthening cross-modal understanding by enhancing video-text expressions in a self-learning manner, which avoids extra workload on data acquisition and prompt design. A CSI module that gradually refines query prompts is designed to generate high-quality captions with rich and diverse expressions, so as to expand the receptive field for cross-modal learning. In addition, multiple learnable experts are introduced as a filter to automatically select the appropriate perspectives of expressions for video-text matching. Our method is non-empirical and self-adaptive in comparison with current text rewriting methods. Experimental results demonstrate that our method achieves SOTA performance on benchmark datasets such as MSR-VTT, MSVD, and DiDemo. Looking ahead, we aspire to explore the variants adaptive to various cross-modal scenarios, thereby aiding a multitude of video-text-related tasks.

### v3

@@ IV-A Datasets
IV-B Implementation Details
@@
Following the standard text preprocessing [23], we split the textual data, including original texts and video-derived captions, into word tokens with a max length of 70 using a CLIP tokenizer. For videos, we sample 8 frames from each item as visual data.
@@
This paper reports the results of our method on multiple backbones, i.e., ViT-B/16, ViT-L/14, ViT-H/14, ViT-G/14, compared to SOTA video-text retrieval methods, showing the effect of model scale on performance as well. Ten captions are derived for each video. Sixteen experts are learned in the ECS module, two of which are activated to select appropriate representations in both training and inference processes. In the training phase, the CSI module is pre-learned to acquire the optimal prompt for caption generation, while the ECS module is co-trained alongside the backbone parameters. In the evaluation phase, multiple captions are derived from each video using the optimal prompt, together with the video itself as video-side expressions. The Captioner and the Prompt engineer are implemented by GPT-4o, a large cross-modal inference model recently released by OpenAI in 2024. The initial prompt asking the Captioner is set as “Generate 10 captions from different perspectives about this video.” The semantic embedding extractor is Equation (1) is implemented by M3-embedding [38]. In addition to the comparison experiment, we encapsulate ExCae as a plug-in unit and test its effect when introduced into existing video-text retrieval methods; specifically, we enhance video expression via CSI and professionalize the video-text matching process by ECS. Ablation studies and analytical experiments are conducted on MSR-VTT using ViT-G/14 as the base model.
@@
Our model is trained on 8 NVIDIA A800 GPUs in a batch size of 16. The pretrained weights from CLIP [23] are used to initialize the text and video encoder in ExCae. The model is trained using the Adamax optimizer with a learning rate of 4e-6. Experimental results are evaluated by standard video-text retrieval metrics: Recall at Rank K (R@K) and Mean Rank (MeanR). R@K denotes the percentage of correct samples in the top K retrieved samples. In this paper, K is set to 1, 5 and 10, respectively, the same as in previous works [4, 33]. MeanR records the average rank of correct items in the retrieved list. The higher value in R@K means better retrieval performance. In contrast, a lower MeanR is preferred in VTR.
IV-C Comparision to State-of-the-arts
@@
The results of our method are compared to those of the SOTAs, including pre-trained foundation models like InternVideo [22], UMT-L [1], mPLUG-2 [32], VAST [2], and ViT-based methods that adapt the image-text models to the video-text task. Among the ViT-based methods, CLIP2TV [7], DRL [27], Clip4Clip [3] and CLIP-VIP [4] require the acquisition of auxiliary data for post-pretraining and/or adaptation. In contrast, TS-Net [24], X-CLIP [9], DMAE [33], Cap4Video++ [15] and TeachClip [26] are learned without extra data.
@@
Results in Table I show that, on the MSR-VTT dataset, our method is superior to all existing methods, including pre-training foundation models and the ViT-based methods. The results of existing methods show that the Top-1 recall accuracy could hardly transcend 50% without assistance from extra data. Instead, great improvement (up to 10%) is obtained when additional data is involved, either they are used for pre-training or fine-tuning. Despite not using any extra data, the Top-1 recall accuracy of our method (ExCae) boosts to 55.0% and 56.1% in text-to-video and video-to-text retrieval, respectively, when using ViT-B/16 as the base model. We are also encouraged by the observation that ExCae’s performance demonstrates an increase in correlation with the enlarged scale of the backbone model, resulting in an around 7% rise of R@1 from ViT-L/14 to Vit-G/14. Our results have comprehensively surpassed the current SOTA pre-training models, such as mPLUG-2 and VAST, when using the Vit-G/14 as the backbone. The rising trend of performance indicates the potential for further improvement of our method through scaling up the base model. Similar results are obtained in experiments on MSVD and DiDeMo datasets, as recorded in Table II. Our results are comparable to or even superior to those of pre-training models. Compared to existing ViT-based methods, ExCae shows a resounding victory, with Top-1 recall accuracy suppressing the best one by 7% on MSVD and 13% on DiDeMo. These results validate that ExCae is robust to VRT tasks for different data sources.
@@
Moreover, we demonstrate several VTR cases to show our results more intuitively. Fig. 4 presents the top three videos retrieved when given a specific caption. It is shown that our method successfully picks the ground-truth videos in the first place. We appreciate seeing that the second and third hit videos are also highly related to the sentences in terms of content. In Fig. 5, we display the text located at the top of the retrieval list when providing a video. Our results are compared to the “Base” obtained without any caption enhancement. The ground-truth texts normally take the first rank when using our method. Compared to the base results, our method preferentially selects samples with semantics and descriptive perspectives similar to the ground truths. For example, our method successfully recognizes keywords of “voices”, “critic”, and “guitar” that are missed by the Base model. This is owing to the ECS module that adaptively learns the cross-modal matching for specific samples, benefiting the recognition of key semantics during video-text retrieval.
@@
IV-D Plug-in Experiment
ExCae is encapsulated and supplemented with existing video-text retrieval methods, whose results are compared to those of the original. As shown in Table III, our method successfully increases retrieval results of different video-text retrieval methods. The Top-1 recall accuracy is boosted by 4.8%, 9.4% and 7.7% for DRL, TSR-Net and Clip4Clip, respectively. It is observed that the results of plug-in experiments are similar to those obtained by our method with the same backbone (refer to Table I). This indicates that ExCae has strong adaptability and expandability, which can flexibly combine with different models and constantly upgrade its performance as the technology evolves.
IV-E Ablation Studies
Our method expands the expression on the video side; thus, this section respectively evaluates the key components, CSI and ECS modules, under different input data cases. To test the effectiveness of the CSI module, we replace the optimal prompt with the initial one to simulate the ablation of the CSI module. The removal of the ECS module is realized by directly encoding all the video-side data as visual embeddings. As shown in Table IV, the R@1 is around 50% when only using the original video as video-side data without adding any components; this result is close to the existing ViT-based methods. The ECS module shows only slight positive effects on the original video. Similar results can be found if we merely change the video-side data from the original videos to video-derived captions. An obvious performance boost occurs when the original videos and video-derived captions are input in combination, validating the indispensability of these two data. In addition, we find that both CSI and ECS modules facilitate cross-modal retrieval. Despite in the absence of the other one, our method achieves over 6% gain in Top-1 recall accuracy. The effectiveness of the CSI and ECS modules proves the superiority of captions’ auto-enhancement and multi-perspective cross-modal matching. The best results are obtained by our full method, where both the CSI and ECS modules play mutually reinforcing roles in promoting video-text understanding.
IV-F Analytical Experiments
TABLE III: Results of plug-in experiments on MSR-VTT dataset.
@@ IV-F Analytical Experiments
Effect of video-derived caption number ($N_{c}$)
Fig. 6 plots the video-text retrieval results when raising $N_{c}$ from 1 to 10. The performance is really poor if only one caption is derived, with a Top-1 recall accuracy of lower than 30% on text-to-video retrieval. This result is inferior to any existing methods, reflecting that insufficient caption enhancement would even damage cross-modal learning. The reason is that all generated captions are forcibly activated to match the text if too few caption samples are input to the ECS module. Consequently, the original text may align with a description from a completely different perspective. Results are much better when increasing the number of generated captions to 3, achieving good performance close to the existing pre-training models. The performance further improves with the increase of the caption number, and the upward trend becomes stable when $N_{c}$ is greater than 7. This experiment suggests an approximate range of $[7,10]$ for the selection of video-derived caption numbers.
Effect of activated expert number ($N_{e}$)
Video-text retrieval results with different $N_{e}$ values are displayed in Fig. 7. We also show the results obtained without the ECS module, labeled as 0 activated expert, as a reference. The retrieval results gradually improve as $N_{e}$ increases from 0 to 2. However, no further improvement in performance is achieved with the subsequent increase of activated experts. This may be because ECS would automatically favor partial experts and suppress the attention to less relevant expressions, although the majority of experts are activated.
TABLE IV: Ablation studies of critical components
@@
V Discussion
V-A Superiority
ExCae attains remarkable video-text retrieval performance primarily due to its efficient narrowing of the video-text modality gap. As shown in Fig. 8 (a), the representations of video-text pairs are drawn closer to each other through self-learning diffusional data augmentation (CSI) and self-adaptive cross-modal matching (ECS). The modality gap $||\Delta||$ [39] is reduced from 0.82 to 0.75, from 0.79 to 0.67, and from 0.84 to 0.74 on MSR-VTT, MSVD and DiDeMo, respectively. Moreover, ExCae has advantages in objectivity, generalization and self-adaptation. 1) ExCae is a data-driven method, in which the query prompts and caption selection are learned according to the training data. No additional human work (despite initialization), such as prompt design and lexicon establishment, is involved in the whole learning process. Thus, ExCae effectively avoids the adverse effects of empiricism. 2) As the Captioner is asked to generate multi-angle captions using a diversity constraint, data with more expression perspectives are provided for model learning. Therefore, ExCae inherits the strength of generalization improvement from data augmentation. 3) ExCae also has personalized ability owing to the parametric learning of the ECS module. The router in the ECS module would adaptively select the best expressions for cross-modal retrieval, thus maintaining the robustness of ExCae in different applications.
@@ V-A Superiority
V-B Convergence
Since the generated captions are changeless once the optimal prompt is determined, the cross-modal post-training of our method owns the same convergence as existing ViT-based methods. Regarding learning the CSI module, the LLM would optimize query prompts in the given direction. With the increase of iteration, the semantic changes in query prompts become smaller, and thus the caption evaluation scores tend to be stable (See Fig. 8 (b)). According to expertise from experiments, the CSI module tends to converge in 400 iterations.
V-C Storage and Efficiency
Although a prompt optimization process is additionally involved, we believe that the storage and computation cost of ExCae is appreciable, especially compared to the pre-training methods. With regard to storage, no extra data or related lexicon needs to be stored. We only need to occupy a very small amount of storage space to record several generated captions in each epoch, which greatly reduces the storage cost. For the computation cost, our method is more efficient than the multi-modal foundation models without a pre-training process. In comparison to the ViT-based methods, the only additional computation is from the training of the CSI module. However, prompt optimization is a one-off training process, and we have also found that the learned prompt has generalization ability across different data sets. Therefore, the efficiency of our approach is comparable to existing VTR methods.
VI Conclusion
This paper proposes strengthening cross-modal understanding by enhancing video expressions rather than text rewriting, which effectively reduces the modality gap without additional workload on data acquisition. We innovatively developed a CSI module that automatically refines query prompts to generate high-quality video captions with rich and diverse expressions. The visual domain is expanded to meet the expression modes and semantic information of texts, so as to facilitate cross-modal understanding. In addition, multiple learnable experts are introduced as a filter to automatically select the appropriate perspectives of expressions for video-text matching. Our method is non-empirical and self-adaptive in comparison with current text rewriting methods. Experimental results demonstrate that our method achieves SOTA performance on benchmark datasets such as MSR-VTT, MSVD, and DiDemo. Looking ahead, we aspire to explore the variants adaptive to various cross-modal scenarios, thereby aiding a multitude of video-text-related tasks.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 62472105, in part by the Natural Science Foundation of Guangdong Province under Grants 2024A1515010186 and 2025A1515011385, as well as the Regional Joint Fund Project of the Basic and Applied Basic Research Foundation of Guangdong Province under Grant 2022A1515140096.